4. $1,000,000 after 40 years with an APR of 10% compounded monthly
Answer #1

PV =1,000,000
Rate per month =10%/12
Number of months =40*12 =480
FV =PV*(1+APR/12)^(40*12) =1000000*(1+10%/12)^480 =53700663.17